OpenStudy (anonymous):

how to solve for pi in SA=2*pi*r^2+2*pi*r*h

OpenStudy (danjs):

you want to solve that for pi ?

OpenStudy (anonymous):

yeah its a manipulation of equations

OpenStudy (danjs):

ok....

OpenStudy (danjs):

really all you have to do is factor the pi out of both terms on the right side, then divide the SA by all that stuff that is left

OpenStudy (anonymous):

but how do i factor out pi on both sides

OpenStudy (danjs):

\[SA = \pi *[~2r^2 + 2rh~]\]

OpenStudy (anonymous):

ohh

OpenStudy (danjs):

yeah , then divide both sides by all tha tin the bra kets

OpenStudy (anonymous):

ok thank you

OpenStudy (danjs):

\[\pi = \frac{ SA }{ 2r^2+2rh } \] for right circular cylinder... welcome